OpenStudy (anonymous):

5x + 7y = 19 I'm having problems finding three solutions, So far I found (1,2). Do I need to just keep on substituting numbers till I get a correct one? It's taking me ages, any faster methods?

jimthompson5910 (jim_thompson5910):

Do you mean integer solutions?

OpenStudy (anonymous):

(-6,7) (1,2) and (8,-3)

OpenStudy (anonymous):

I take it you just kept on substituting numbers till you got a valid solution huh :/ Damn was hoping for there to be some other way to do this faster, thanks for the help.

OpenStudy (anonymous):

yeah well if u got more just lemme know

jimthompson5910 (jim_thompson5910):

if you know one integer solution for Ax+By = C, call it (x1, y1), then you can find any integer solution using the equations x = x1 + B*k y = x2 - A*k for any integer k For example, if k = 2, then x = x1 + B*k y = x2 - A*k x = 1 + 7*k y = 2 - 5*k x = 1 + 7*2 y = 2 - 5*2 x = 15 y = -8 So (15,-8) is another integer solution